### v1

@@ Abstract
The complex systems keyword diagram generated by the author in 2010 has been used widely in a variety of educational and outreach purposes, but it definitely needs a major update and reorganization. This short paper reports our recent attempt to update the keyword diagram using information collected from the following multiple sources: (a) collective feedback posted on social media, (b) recent reference books on complex systems and network science, (c) online resources on complex systems, and (d) keyword search hits obtained using OpenAlex, an open-access bibliographic catalogue of scientific publications. The data (a), (b) and (c) were used to incorporate the research community’s internal perceptions of the relevant topics, whereas the data (d) was used to obtain more objective measurements of the keywords’ relevance and associations from publications made in complex systems science. Results revealed differences and overlaps between public perception and actual usage of keywords in publications on complex systems. Four topical communities were obtained from the keyword association network, although they were highly intertwined with each other. We hope that the resulting network visualization of complex systems keywords provides a more up-to-date, accurate topic map of the field of complex systems as of today.
@@ 2 Data and Methods
The data (a), (b) and (c) were used to incorporate the research community’s internal perceptions of the relevant topics, whereas the data (d) was used to obtain more objective measurements of the keywords’ relevance and associations from publications made in complex systems science. In (a)-(c), the relevant keywords were manually selected and curated by the author. In (d), the overall relevance of each keyword to complex systems was characterized by its total search hit multiplied by its visibility boost [15] when it was searched together with “complex systems” OR “complexity science”. Namely, if the relevance score of a keyword is high, that implies that the keyword was more manifested in the context of complex systems or complexity science (see [15] for details).
One problem faced in collecting the data (d) above was that many combinations of keywords resulted in superficial large search hits that might not correctly capture their actual scientific relatedness, especially when both of them were common words/phrases. To avoid such misleading measurements and improve the quality of network reconstruction, the entire relevant keyword list and possible associations of the pairs of keywords were manually reviewed and curated by the author from the viewpoint of complex systems science and engineering. The keyword association network was then constructed by assigning edge weights between a pair of associated keywords according to the OpenAlex search hit of the pair multiplied by both keywords’ relevance scores.
@@ 3 Results
However, note that these community labels are no more than just to aid the visibility and organization of the diagram, and they are in fact heavily intertwined with each other if the entire keyword association network is visualized using a conventional network layout algorithm (Fig. 4). In this entire network visualization, the keywords that form the conceptual core of complex systems naturally come at the center, such as “complexity”, “emergence”, “self-organization”, “information”, “structure”, and “patterns”, even though some of them may not have gained a high relevance score in the literature. The broad range of different topical areas spread from the central core, covering biology, engineering, mathematics, physics, philosophy and others. Readers are encouraged to zoom into this diagram to explore the details and how the topics are placed relative to others in this new keyword diagram.
@@ 4 Conclusions
In this study, the keyword diagram of complex systems has been updated in the form of a complex weighted network of associated keywords by using the collective feedback from the community/public and the data obtained from the OpenAlex scientific literature database. The comparison between the data obtained from these two data sources highlighted several topical differences, indicating that there might be some gaps between public perception and actual usage of keywords in scientific publications on complex systems. However, the visualization of the entire association network of keywords did place frequently mentioned keywords near the center of the network (despite their relatively lower relevance scores), implying that those central keywords do play an important role in connecting various other keywords together even in the scientific publications.
The attempt to find topical clusters using the Louvain method produced four communities, but they were all heavily intermingled with each other, resulting in a more nuanced, networked view of the complex systems keywords. Nonetheless, these communities helped recognize the large-scale topical dimensions in the generated keyword diagram that consist of the theory-application axis and the computational-analytical axis. It is our hope that this networked keyword diagram represents the actual state of the field of complex systems science and the breadth and depth of its topics more accurately than the original diagram that might have been overly simplistic and hierarchical.
There are still many limitations in this work. One critical problem is that it still used manually curated keyword associations that would have been heavily influenced by the author’s subjective evaluation and filtering. It remains an open question how to select scientifically meaningful associations between keywords solely from the literature data without manual curation. Another problem is the lack of comparison of the obtained result with results of other similar recent efforts of complex systems topic mapping (e.g., [18, 19]). It would be interesting and useful to integrate the results of all of those related projects to build a meta-level keyword map. Finally, exploration of a complex topic network would be extremely difficult on a static visualization and hence it may not carry much pedagogical values. It would be much more valuable and useful to have an interactive, dynamic interface by which users can actively explore the relevant keywords, which is currently under development.

### v2

@@ Abstract
The complex systems keyword diagram generated by the author in 2010 has been used widely in a variety of educational and outreach purposes, but it definitely needs a major update and reorganization. This short paper reports our recent attempt to update the keyword diagram using information collected from the following multiple sources: (a) collective feedback posted on social media, (b) recent reference books on complex systems and network science, (c) online resources on complex systems, and (d) keyword search hits obtained using OpenAlex, an open-access bibliographic catalogue of scientific publications. The data (a), (b) and (c) were used to incorporate the research community’s and other public communities’ perceptions of the relevant topics, whereas the data (d) was used to obtain more objective measurements of the keywords’ relevance and associations from publications made in complex systems science. Results revealed differences and overlaps between public perception and actual usage of keywords in publications on complex systems. Four topical communities were obtained from the keyword association network, although they were highly intertwined with each other. We hope that the resulting network visualization of complex systems keywords provides a more up-to-date, accurate topic map of the field of complex systems as of today.
@@ 2 Data and Methods
The data (a), (b) and (c) were used to incorporate the research community’s and other public communities’ perceptions of the relevant topics, whereas the data (d) was used to obtain more objective measurements of the keywords’ relevance and associations from publications made in complex systems science. In (a)-(c), the relevant keywords were manually selected and curated by the author. In (d), the overall relevance of each keyword to complex systems was characterized by its total search hit multiplied by its visibility boost [15] when it was searched together with “complex systems” OR “complexity science”. Namely, if the relevance score of a keyword is high, that implies that the keyword was more manifested in the context of complex systems or complexity science (see [15] for details).
One problem faced in collecting the data (d) above was that many combinations of keywords resulted in superficial large search hits that might not correctly capture their actual scientific relatedness, especially when both of them were common words/phrases. To avoid such misleading measurements and improve the quality of network reconstruction, the entire relevant keyword list and possible associations of the pairs of keywords were manually reviewed and curated by the author from the viewpoint of complex systems science and engineering. The keyword association network was then constructed by assigning edge weights between a pair of associated keywords according to the OpenAlex search hit of the pair (searched together with “complex systems” OR “complexity science”) multiplied by both keywords’ relevance scores.
@@ 3 Results
However, note that these community labels are no more than just to aid the visibility and organization of the diagram, and they are in fact heavily intertwined with each other if the entire keyword association network is visualized using a conventional network layout algorithm, which does not show any immediately visible community structure (Fig. 4). In this entire network visualization, the keywords that form the conceptual core of complex systems naturally come at the center, such as “complexity”, “emergence”, “self-organization”, “information”, “structure”, and “patterns”, even though some of them may not have gained a high relevance score in the literature. The broad range of different topical areas spread from the central core, covering biology, engineering, mathematics, physics, philosophy and others. Readers are encouraged to zoom into this diagram to explore the details and how the topics are placed relative to others in this new keyword diagram.
@@ 4 Conclusions
In this study, the keyword diagram of complex systems has been updated in the form of a complex weighted network of associated keywords by using the collective feedback from the community/public and the data obtained from the OpenAlex scientific literature database. The comparison between the data obtained from these two data sources highlighted several topical differences, indicating that there might be some gaps between public perception and actual usage of keywords in scientific publications on complex systems. Such comparison of word uses between public discourse and scientific publications can be a general research methodology potentially useful for other similar topic mapping studies too. Meanwhile, the visualization of the entire association network of keywords did place frequently mentioned keywords near the center of the network (despite their relatively lower relevance scores), implying that those central keywords do play an important role in connecting various other keywords together even in the scientific publications.
The attempt to find topical clusters using the Louvain method produced four communities, but they were all heavily intermingled with each other, resulting in a more nuanced, networked view of the complex systems keywords. Nonetheless, these results of community detection helped recognize the large-scale topical dimensions in the generated keyword diagram that consist of the theory-application axis and the computational-analytical axis. It is our hope that this networked keyword diagram represents the actual state of the field of complex systems science and the breadth and depth of its topics more accurately than the original diagram that might have been overly simplistic and hierarchical.
There are still many limitations and possible future directions in this work. One critical problem is that it still used manually curated keyword associations that would have been heavily influenced by the author’s subjective evaluation and filtering. It remains an open question how to select scientifically meaningful associations between keywords solely from the literature data without manual curation. Another major problem is that the two topical axes proposed in this study were simply based on crude interpretation of a network visualization that was made using a specific heuristic network layout algorithm. It would be interesting and more convincing to see if similar axes could be extracted more analytically using a more rigorous dimensionality reduction method. Furthermore, yet another problem is the lack of comparison of the obtained result with results of other similar recent efforts of complex systems topic mapping (e.g., [18, 19]). It would be interesting and useful to integrate the results of all of those related projects to build a meta-level keyword map, hopefully as a community effort coordinated on a regular basis so that we can track how the relevant topics in complex systems change over time. Finally, exploration of a complex topic network would be extremely difficult on a static visualization and hence it may not carry the same level of pedagogical values as the original 2010 diagram had. It would be much more valuable and useful to have an interactive, dynamic interface by which users can actively explore the relevant keywords and organize the topics into categories in their preferred ways, which is currently under development.